If 2xy=0 and x/2 = 3/4, which of the following must be true?
1. 2y=-3
2. x=-12y
3. 2x=3
A) none
B) 1. only
C) 2. only
D) 3. only
E) 1. and 2. only

Answer by stanbon(75887)

If 2xy=0 and x/2 = 3/4, which of the following must be true?
-----
Note:: x = 2(3/4) = 3/2 ; then 2y(3/2) = 0, so y = 0
----------------------------------
1. 2y=-3
2. x=-12y
3. 2x=3
A) none
B) 1. only
C) 2. only
D) 3. only
E) 1. and 2. only
=======
Answer:: D
-----------------
Cheers,
Stan H.
-----------------